Find lim x tends to 0 1-cos2x /x2sinx2


1
Expert's answer
2021-06-11T02:47:02-0400
"\lim\limits_{x\to 0}\dfrac{1-\cos^2(x)}{x^2\sin(x^2)}=\lim\limits_{x\to 0}\dfrac{\sin^2(x)}{x^2\sin(x^2)}"

"=\lim\limits_{x\to 0}\bigg((\dfrac{\sin(x)}{x})^2\cdot\dfrac{1}{\sin(x^2)}\bigg)"

"=(1)^2\lim\limits_{x\to 0}(\dfrac{1}{\sin(x^2)})=\infin"